OpenStudy (anonymous):

MEDAL AND FAN LAST QUESTION!!! Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6

OpenStudy (anonymous):

A) y+3=-6(x+2) B) y+2=6(x+3) C) y-3=-6(x-2) D) y-2=6(x-3)

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

Do you know this one?

OpenStudy (anonymous):

I think its D but I'm not sure

OpenStudy (anonymous):

Wait it can't be D...

OpenStudy (anonymous):

y=-3 has to be right?

OpenStudy (anonymous):

Since the y is -3...

OpenStudy (anonymous):

\[y - y_1 = m(x - x_1)\] x1 = -2 y1 = -3 m = -6 Put these values

OpenStudy (anonymous):

but how would I find the second y and the second x if there is none...

OpenStudy (anonymous):

The slope is -6 and the points are (-2,-3) so that would mean we would have to find the equation that passes through the point.... Wouldn't that mean it has to pass through either -2, or -3?

OpenStudy (anonymous):

@Ogzii can you help me?

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

Put value of x1, y1, m in above equation and nothing else. you will have your ans

OpenStudy (anonymous):

WAIT :D

OpenStudy (anonymous):

y- (-3) =6 (x--2) So that would be y+3=6(x+2) :D

OpenStudy (anonymous):

am I right @sangya21

OpenStudy (anonymous):

y+3=-6(x+2)?

OpenStudy (anonymous):

yes no?

OpenStudy (anonymous):

Coreect

OpenStudy (anonymous):

YAY thanksies :D